OpenStudy (anonymous):

Please help Find an exact value. sine of negative eleven pi divided by twelve.

OpenStudy (anonymous):

So it's sin (-11pi/12)

OpenStudy (anonymous):

-11pi/12 is half of -11pi/6 so you can use a half angle formula. \[\sin \frac{ \theta }{ 2 }=\pm \sqrt{\frac{ 1-\cos \theta }{ 2 }}\] where Θ=-pi/6 and you choose + or - based on the quadrant

OpenStudy (anonymous):

sorry Θ=-11pi/6

OpenStudy (anonymous):

@peachpi My options are quantity square root of six plus square root of two divided by four. quantity negative square root of six minus square root of two divided by four. quantity square root of two minus square root of six divided by four. quantity square root of six minus square root of two divided by four.

OpenStudy (anonymous):

I wouldn't know which one of those it is because I didn't solve the problem. All you have to do is plug in -11pi/6.

OpenStudy (anonymous):

Thank you!